# Scene-Adaptive Motion Planning with Explicit Mixture of Experts and Interaction-Oriented Optimization

### v2

@@ III-E2 Interaction-oriented prediciton
The discourse in Section 2 reveals that prevalent models typically overlook the dynamic interplay between the ego vehicle and other vehicles in trajectory prediction. Recognizing the real-world implications of such interactions, this study integrates the interaction of the ego vehicle’s planned trajectory with those of other vehicles into the prediction framework. As illustrated in Figure 1, this is achieved by introducing a Pred Decoder module. This module synthesizes the agent encoding from the Scene Encoder with the ego vehicle trajectory features from the MOE Decoder via a cross-attention mechanism, facilitating interaction.
@@ IV-A Experimental Setup
The training and evaluation of the model were conducted using the NuPlan [50] dataset. From this dataset, 3 million samples, spread across various scenarios, constituted the training sample pool. During preprocessing, samples were categorized into seven scenario types as outlined in Section 3.2. However, not all samples were utilized during the training process. Instead, the size of the training dataset was regulated by controlling the number of each scenario, detailed in Section 4.2.
@@
The model was evaluated using closed-loop non-reactive (NR) simulations with Val14 [7], Test14, and Test14-hard [5] benchmarks. Although Nuplan offers reactive simulations, these are not considered robust due to their simplistic rule-based interactions. The industry uses non-reactive obstacle-based interactions and assigns the main responsibility for collisions, as done in Nuplan’s NR simulations. Therefore, we only use NR as the standard for evaluation. Detailed evaluation metrics can be found in [51].
@@
Name Description Values
lr learning rate 1e-3
$T_{f}$ Future Steps 80
$T_{h}$ History Steps 20
$L_{e}$ Num. scene encoder layers 4
$L_{d}$ Num. decoder layers 4
$N_{E}$ Num. experts 7
$K_{a}$ Num. anchors 24
$k_{R}$ Weight decay factor 0.2
$D$ Dimention of features 128
$D_{h}$ Hidden dim of expert 512
@@ IV-B Results
DiffusionPlanner 89.87 97.03 94.40 98.74 96.89 91.20 98.11
EMOE-Planner(Ours) 93.06 97.98 94.77 99.54 98.16 92.35 97.87
@@
According to the results in Table II, our model outperforms all other learning-based methods across all testing scenarios, particularly excelling in the val14 and test14-hard scenarios where it exceeds the state-of-the-art (SOTA) results by a minimum of 3 points. In the val14 scenarios, the performance of the EMOE-Planner closely approaches that of expert scores.Among rule-based methods, when implementing the same post-processing technique used in Pluto and DiffusionPlanner with our EMOE-Planner, it achieves the highest performance in the val14 and test14-hard scenarios, albeit slightly lower than DiffusionPlanner in the test14 scenario. It is noteworthy that the performance of EMOE-Planner with and without post-processing shows slight differences, suggesting that the trajectories produced by the model are of high quality, potentially diminishing the necessity for post-processing, aligning with our objectives. Additionally, the EMoE-Planner, without post-processing, scores only 0.2 points lower than PDM-Closed, which is a rule-based but state-of-the-art algorithm in the NuPlan Challenge, on the val14 test set. On all other datasets, it outperforms PDM-Closed. To our knowledge, we are the first pure learning model to achieve superior performance in closed-loop simulations compared to rule-based algorithms on the NuPlan dataset. In comparison, only when post-processing is applied do the Pluto and Diffusion Planners achieve scores higher than those of PDM-Closed. Furthermore, on the Test14-hard dataset, EMOE-Planner significantly outperforms PDM-Closed. This indicates that rule-based algorithms struggle to handle complex scenarios, while our model continues to perform well in these situations.
As indicated in Table III, our model generally outperforms both Pluto and DiffusionPlanner in all metrics except speed. Notably, it excels in comfort and progress metrics. The low progress score of Pluto aligns with observations from simulations where Pluto’s trajectories tend to be conservative, often slower than expert trajectories. In the appendix, we provide detailed trajectories for various scenarios to give a comprehensive understanding of the situations discussed here. Our loss function accounts for vehicle interactions, which effectively optimizes the trajectory progress, particularly in overtaking scenarios, thus avoiding excessive conservatism while preventing collisions.
@@
In the EMOE architecture, the distribution of scenario quantities in the training data is crucial for model performance. The second row in Table IV details the scenario counts in the training dataset. ST-J, LT-J, RT-J means go straight, turn left and turn right in junctions. ST means go straight in non-junction areas. UT is u-turn and RA is roundabout. During training, we capped each scenario type at 200,000 from a pool of 3,000,000 samples as discussed in section 4.1. The first four scenario types reached this cap, while the last three did not due to insufficient data in the 3M training database, resulting in a total of 1,083,804 samples used for training. We categorized the scenarios in val14 using the method outlined in the appendix and conducted separate simulation assessments for each scenario type, with results presented in Table IV.
@@ IV-C Ablation Studies
This section presents ablation studies on three components: EMOE, scene-specific query(SSQ), and interaction-oriented loss and prediction(I-Loss), using the val14 dataset. Table 6 displays the scores following the removal of each module.
@@
Version Score Collisions TTC Drivable Comfort Progress Speed AvgADE AvgFDE
Base 93.06 97.98 94.77 99.54 98.16 92.35 97.87 0.71 2.12
w/o EMOE 92.33 97.43 94.49 99.54 98.89 92.26 97.84 0.71 2.12
w/o SSQ 92.51 97.99 95.41 99.17 97.06 91.72 97.92 0.71 2.13
w/o I-Loss 91.44 96.78 93.85 99.72 97.88 91.88 97.89 0.73 2.20
Ablation experiments reveal that interaction-oriented loss and prediction significantly enhance performance, underscoring the importance of interaction. EMOE and scene-specific query also play crucial roles in boosting model performance and these two are used in conjunction. The scene-specific query guides the model’s focus to key scene areas, and EMOE processes that area’s information with a scene-adapted expert. Additionally, Table V shows that factoring in interactions yields better predictions for surrounding agents.
IV-D Simulation Results
Figures 4 through 7 illustrate the closed-loop simulation results for four distinct scenarios. We present snapshots at 0s, 5s, 10s, and 15s for both Pluto and EMOE-Planner. In Figure  4(b), it is clear that Pluto collides with a stationary vehicle during a right turn, while EMOE-Planner navigates a safe trajectory. The trajectory by EMOE-Planner in Figure 4(f) initially aligns more closely with the ground truth, attributable to the incorporation of temporal-weighted regression loss which prioritizes near-term planning and enhances learning from the expert’s obstacle avoidance maneuvers. Conversely, Pluto prioritizes overall trajectory resemblance to the expert, potentially compromising local accuracy for global alignment, leading to less optimal near-term planning.
Figure 4: Closed-loop simulation in scenario1. Subfigures (a) through (d) depict snapshots at 0s, 5s, 10s, and 15s for Pluto, while subfigures (e) through (h) represent snapshots at corresponding times for EMOE-Planner.
In the scenario depicted in Figure  5, Pluto halts at the intersection instead of executing the expected right turn, while EMOE-Planner successfully completes it. The reason for Pluto’s halt is difficult to ascertain; however, it is evident that the expert network in EMOE-Planner more effectively manages such irregular scenarios.
Figure 5: Closed-loop simulation in scenario2.
Figure 6 depicts a right turn at a busy intersection, where both Pluto and EMOE-Planner execute the maneuver correctly. However, Figure 6(b) shows that Pluto adopts a notably cautious approach, trailing almost one car length behind the expert at 5 seconds, and by Figures 6(d), it falls approximately 15 meters behind by 15 seconds. In contrast, EMOE-Planner is only about 7 meters behind in Figure 6(h). Pluto’s trajectory does not account for interactions with other vehicles, leading to a safer but more conservative path, while EMOE-Planner’s interaction-aware planning yields a trajectory closer to that of the expert.
Figure 6: Closed-loop simulation in scenario3.
In Figure 7, the scenario involves driving straight through an intersection. Here, Pluto unnecessarily changes lanes mid-intersection likely due to a perceived collision threat from the vehicle in the red box in Figure 7(b), prompting a lane change that is reversed in Figure 7(c). EMOE-Planner, however, evaluates the collision potential more accurately, maintaining a straight trajectory and successfully navigating the intersection. However, although EMOE-Planner proceeds straight, the trajectory remains conservative due to the high traffic in the scenario. The current framework does not yet consider the prediction of other vehicles when planning the ego vehicle’s trajectory, which will be a focus of our future work.
Figure 7: Closed-loop simulation in scenario4.

### v3

@@ III-E2 Interaction-oriented prediciton
The discourse in Section II reveals that prevalent models typically overlook the dynamic interplay between the ego vehicle and other vehicles in trajectory prediction. Recognizing the real-world implications of such interactions, this study integrates the interaction of the ego vehicle’s planned trajectory with those of other vehicles into the prediction framework. As illustrated in Figure 1, this is achieved by introducing a Pred Decoder module. This module synthesizes the agent encoding from the Scene Encoder with the ego vehicle trajectory features from the MoE Decoder via a cross-attention mechanism, facilitating interaction.
@@ IV-A Experimental Setup
The training and evaluation of the model were conducted using the NuPlan [3] dataset. From this dataset, 3 million samples, spread across various scenarios, constituted the training sample pool. During preprocessing, samples were categorized into seven scenario types as outlined in Section III-D. However, not all samples were utilized during the training process. Instead, the size of the training dataset was regulated by controlling the number of each scenario, detailed in Section IV-B.
@@
The model was evaluated using closed-loop non-reactive (NR) simulations with Val14 [10], Test14, and Test14-hard [7] benchmarks. Although Nuplan offers reactive simulations, these are not considered robust due to their simplistic rule-based interactions. The industry uses non-reactive obstacle-based interactions and assigns the main responsibility for collisions, as done in Nuplan’s NR simulations. Therefore, we only use NR as the standard for evaluation. The evaluation metrics are the same with those in [25] and [30]. For a fair and consistent comparison with existing works, all evaluation scores are scaled by 100 compared to official scores. Detailed information is in the supplementary materails.
@@
  • GameFormer[20] – The model enhances prediction accuracy for both the ego and surrounding vehicles by modeling their interactions based on game theory.
@@
Name Description Values Name Description Values
lr learning rate 1e-3 $L_{e}$ Num. scene encoder layers 4
$T_{f}$ Future Steps 80 $k_{R}$ Weight decay factor 0.02
$T_{h}$ History Steps 20 $L_{d}$ Num. decoder layers 4
$N_{E}$ Num. experts 7 $D$ Dimention of features 128
$K_{a}$ Num. anchors 24 $D_{h}$ Hidden dim of expert 512
@@ IV-B Results
DiffusionPlanner w/o post. 89.87 97.03 94.40 98.74 96.89 91.20 98.11
EMoE-Planner w/o post. (Ours) 93.06 97.98 94.77 99.54 98.16 92.35 97.87
@@
According to the results in Table II, our model outperforms all other learning-based methods across all testing scenarios, particularly excelling in the val14 and test14-hard scenarios where it exceeds the state-of-the-art (SOTA) results by a minimum of 3 points. In the val14 scenarios, the performance of the EMoE-Planner closely approaches that of expert scores. Among rule-based methods, when implementing the same post-processing technique used in Pluto and DiffusionPlanner with our EMoE-Planner, it achieves the highest performance in the val14 and test14-hard scenarios, albeit slightly lower than DiffusionPlanner in the test14 scenario. It is noteworthy that the performance of EMoE-Planner with and without post-processing shows slight differences, suggesting that the trajectories produced by the model are of high quality, potentially diminishing the necessity for post-processing, aligning with our objectives. Additionally, the EMoE-Planner, without post-processing, scores only 0.02 points lower than PDM-Closed, which is a rule-based but state-of-the-art algorithm in the NuPlan Challenge, on the val14 test set. On all other datasets, it outperforms PDM-Closed. To our knowledge, we are the first pure learning model to achieve superior performance in closed-loop simulations compared to rule-based algorithms on the NuPlan dataset. In comparison, only when post-processing is applied do the Pluto and Diffusion Planners achieve scores higher than those of PDM-Closed. Furthermore, on the Test14-hard dataset, EMoE-Planner significantly outperforms PDM-Closed. This indicates that rule-based algorithms struggle to handle complex scenarios, while our model continues to perform well in these situations.
As indicated in Table III, our model generally outperforms both Pluto and DiffusionPlanner in all metrics except speed. The EMoE-Planner achieves a higher progress score, indicating superior traffic throughput efficiency, while maintaining a slightly lower speed metric, reflecting a more assertive yet rational driving style. This is balanced by notable improvements in safety-critical metrics: fewer collisions and a higher TTC, demonstrating enhanced safety and better risk anticipation. This suggests that the more assertive speed profile primarily occurs in low-risk scenarios. Overall, the trajectories generated by the EMoE-Planner are both faster and safer, closely resembling human driving performance.
One key reason for this balanced performance is the design of the Interaction-Oriented Loss, which explicitly accounts for dynamic interactions in scenarios like overtaking to generate trajectories that are neither overly conservative nor risky. Another reason is the combined effect of the EMoE architecture and scene-specific queries, which simplifies expert learning per scenario and focuses attention on critical spatial areas implicitly enhancing the overall trajectory quality and leading to an overall improvement in metric scores.
@@
In the EMoE architecture, the distribution of scenario quantities in the training data is crucial for model performance. The second and fifth column in Table IV show the scenario counts in the training dataset. ST-J, LT-J, RT-J means go straight, turn left and turn right in junctions. ST means go straight in non-junction areas. UT is u-turn and RA is roundabout. During training, we capped each scenario type at 200,000 from a pool of 3,000,000 samples as discussed in section 4.1. The first four scenario types reached this cap, while the last three did not due to insufficient data in the 3M training database, resulting in a total of 1,083,804 samples used for training. We conduct separate simulation assessments for each scenario type, with results presented in Table IV.
@@ IV-C Ablation Studies
This section presents ablation studies of each component and the weith decay factor $k_{R}$ with the val14 dataset.
Ablation of each componets
: We conduc ablation studies of three components: EMoE, scene-specific query(SSQ), and interaction-oriented loss and prediction(I-Loss). Table V displays the scores following the removal of each module. Ablation experiments reveal that interaction-oriented loss and prediction significantly enhance performance, underscoring the importance of interaction. EMoE and scene-specific query also play crucial roles in boosting model performance and these two are used in conjunction. The scene-specific query guides the model’s focus to key scene areas, and EMoE processes that area’s information with a scene-adapted expert. Additionally, Table V shows that factoring in interactions yields better predictions for surrounding agents.
@@
Version Score AvgADE AvgFDE
Base 93.06 0.71 2.12
w/o EMoE 92.33 0.71 2.12
w/o SSQ 92.51 0.71 2.13
w/o I-Loss 91.44 0.73 2.20
Weight decay factor $k_{R}$
: According to Equation 11 and  12, $k_{R}$ has a significant impact on the overall regression loss. This set of experiments is conducted to test the influence of this parameter, and the results are presented in Fig 4. According to Fig 4, the optimal performance is achieved when $k_{R}$ is set to 0.02. A larger $k_{R}$ causes the model to over-focus on short-term trajectories, which degrades long-term planning performance. Conversely, decreasing $k_{R}$ results in insufficient attention to short-term interactions, leading to poorer performance in immediate scenarios.
Figure 4: Weight decay factor
IV-D Computation Efficiency
This section presents a comparative analysis of the computational efficiency between EMoE-Planner, Pluto and DiffusionPlanner. The comparison is conducted across three key dimensions: the memory usage, the average inference latency, and the average floating point operations (FLOPs). The detailed results are summarized in Table VI.
TABLE VI: Computation Efficiency.
Planners Memory(M) FLOPs(G) Latency(ms)
Pluto w/o post. 4.24 0.82 20.75
DiffusionPlanner w/o post. 6.04 1.54 77.44
EMoE-Planner w/o post. 8.91 2.42 36.27
The EMoE-Planner incorporates 24 additional experts across its four decoder layers compared to other models, resulting in the largest memory usage. Furthermore, due to the optimized prediction decoder and the integration of scene-specific queries, the EMoE-Planner also exhibits the highest FLOPs. However, its inference latency remains competitive, which is only marginally higher than Pluto as only a single expert is active per inference. From an engineering perspective, the EMoE-Planner fully satisfies the real-time requirements for autonomous driving applications.
IV-E Simulation Results
Figures 5 and 6 illustrate two of the least common scenario types among the seven defined categories: roundabout and u-turn maneuvers. We present snapshots at key time points for DiffusionPlanner, Pluto and EMoE-Planner.
Figure 5: Closed-loop simulation of roundabout. Subfigures (a) through (d) depict snapshots at 0s, 8s, 12s, and 15s for DiffusionPlanner, subfigures (e) through (h) represent snapshots at corresponding times for Pluto and subfigures(i) through (l) denote snapshots for EMoE-Planner. The ego vehicle is represented by an orange rectangle, while a gray rectangle denotes its ground truth position at the same timestamp. The planned trajectory is shown as a colored solid line, the historical trajectory in purple, and the ground-truth historical trajectory in blue. The candidate planned trajectories are displayed as gray lines.
In Figure 5(b)(f)(j) at 8 seconds, Pluto collides with a left-side vehicle, while DiffusionPlanner maintains an uncomfortably close distance. In contrast, EMoE-Planner keeps a safe gap. By 12 seconds shown in Figure 5(c)(g)(k), DiffusionPlanner collides with a vehicle on its right, and Pluto continues its collision on the left. EMoE-Planner, however, consistently maintains safe distances. Ultimately, only EMoE-Planner navigates the entire scenario without any collision. Comparing positions at 15s (Figure 5(d)(h)(l)), EMoE-Planner travels the farthest, indicating the highest efficiency. This performance stems from two key design features of EMoE-Planner. First, its interaction-oriented loss explicitly accounts for interactions with other vehicles, which leads to significantly better collision avoidance. Second, the specialized expert dedicated to roundabout scenarios within the EMoE architecture allows the model to generate trajectories that are better adapted to the high-curvature geometry of roundabouts. This results in more reasonable and efficient paths compared to other planners.
In Figure 6, all three planners safely navigate the U-turn. However, their positions at 15 seconds reveals that EMoE-Planner travels the farthest, even slightly exceeding the ground-truth trajectory by nearly one car length. Furthermore, according to historical trajectories at 15 seconds, DiffusionPlanner’s path is excessively wide (”outer loop”), whereas EMoE-Planner selects a tighter, more efficient inner path. Similar to the roundabout case, the dedicated U-turn expert within the EMoE architecture enables better adaptation to the scenario’s geometry, resulting in superior trajectories.
This section presents simulation results for two of the rarest and most challenging scenario types, where EMoE-Planner demonstrates significant performance improvements. In the supplementary material, we provide comparative results for the remaining five scenario types, including some particularly difficult cases and scenarios where EMoE-Planner still faces challenges. Across nearly all scenario types, EMoE-Planner outperforms both DiffusionPlanner and Pluto, especially in terms of safety and efficiency.
Figure 6: Closed-loop simulation of u-turn scenario.
@@ V Conculsion
However, EMoE-Planner still has certain limitations. According to the results in Section IV(E) and supplementary material, our model currently underperforms in terms of traffic efficiency and in handling complex interactive scenarios compared to human drivers. Moreover, all evaluations in this work have been conducted within the nuPlan simulation environment. In future work, we aim to further refine the model to improve its performance in complex interactions and enhance its efficiency while maintaining safety. Most importantly, we plan to deploy and validate our model on real autonomous vehicles to advance its practical application.
@@ Acknowledgments
SUPPLEMENTARY MATERIAL
VI-A Evaluation Metrics
The evaluation metrics used in our study are the official benchmarks provided by the nuPlan platform [30], which have been widely adopted by existing research [6, 48] utilizing the nuPlan dataset. The specific metrics are as follows:
  • – No ego at-fault collisions: A collision is defined as the event of ego’s bounding box intersecting another agent’s bounding box. Except for collisions that occur while the ego vehicle is stationary or when an active track hits the ego from the rear, all other collisions are considered at-fault. This metric score increases as the number and severity of at-fault collisions decrease. This metric is used to evaluate the safety of the trajectory.
  • – Drivable area compliance: This metric identifies the frames when ego drives outside the drivable area. This metric is a binary criterion. If any frame within a clip is judged to be outside the drivable area, the score is 0. A score of 1 is only achieved if the trajectory remains within the drivable area for all frames.
  • Time to Collision (TTC) within bound: TTC is defined as the time required for ego and another track to collide if they continue at their present speed and heading. This metric is also boolean, 1 if TTC is higher than a minimum lower bound, for example 0.95s, else 0. TTC is indicative of driving caution, a higher TTC value signifies a more conservative driving style.
  • – Progress: This metric evaluates how much of the expert’s path the ego vehicle has covered. It is calculated by comparing the distance traveled by the ego vehicle to the total length of the expert’s route. A higher value for this metric directly indicates higher throughput efficiency of the planned trajectory.
  • – Speed limit compliance: This metric evaluates if ego’s speed exceeds the associated speed limit in the map. The score for this metric is determined by considering both the duration and the severity of the overspeeding violation.
  • – Comfort: The comfort level is assessed by measuring the extreme values of critical kinematic parameters, encompassing longitudinal and lateral accelerations, yaw rate, yaw acceleration, and jerk.
Given the complexity involved in the detailed calculation of these metrics, the official nuPlan website [30] remains the definitive source for complete information.
VI-B Additional Simulation Results
Figure 711 present closed-loop simulation results for the five additional scenario types beyond those discussed in the main text. For each scenario, we captured screenshots at the critical timestep specific to each model, as their key decision points may differ. The primary challenge varies per scenario: left turns require balancing turning radius to avoid collisions with inner or outer lane vehicles; right turns must manage conflicts with oncoming cross-traffic; non-intersection going straight focuses on avoiding stationary vehicles during lane changes; going straight within intersections necessitates careful acceleration control to prevent rear-end or front-end collisions; and other complex scenarios like merging prioritize correct lane selection and dynamic collision risk management.
In Figure 7, the DiffusionPlanner’s trajectory exhibits an excessively wide outer swing, resulting in a collision with a vehicle in the oncoming lane at 7s. Meanwhile, the PlutoPlanner passes dangerously close to an oncoming vehicle at the 9-second mark and subsequently collides with a red car in the opposing lane at 10 seconds. In contrast, the EMoE-Planner safely navigates through the intersection with a trajectory that closely aligns with the ground-truth path. Both in terms of efficiency and safety, EMoE-Planner performs the best in the right-turn scenario.
Figure 7: Closed-loop simulation of right-turn scenario. Subfigures (a) through (d) depict snapshots at 0s, 5s, 7s, and 10s for DiffusionPlanner, subfigures (e) through (h) represent snapshots at 0s, 5s, 9s, and 10s for Pluto and subfigures(i) through (l) denote snapshots at 0s, 5s, 8s, and 9s for EMoE-Planner.
Figure 8 shows the result in left-turn scenario. The DiffusionPlanner’s trajectory exhibits an excessively wide outer loop as in Figure 7, causing it to collide with a vehicle on the right during its return at the 14-second mark. Meanwhile, the PlutoPlanner suffers from inefficient steering, resulting in a rear-end collision by a following vehicle at 10 seconds. In contrast, the EMoE-Planner generates a trajectory nearly identical to the ground truth, completing the turn both safely and efficiently.
Figures 9 and 10 show the results for straight driving on a non-intersection road and going straight through an intersection, respectively. In the non-intersection scenario, both DiffusionPlanner and Pluto experience collisions at different time steps, while EMoE-Planner completes the maneuver safely. Moreover, based on the ego vehicle’s position at 15s, EMoE-Planner again achieves the highest efficiency. In Figure 10 the intersection straight-driving scenario, DiffusionPlanner and Pluto both exhibit slow starts, leading to inefficient progression and rear-end collisions by following vehicles at 6s and 10s, respectively. In contrast, EMoE‑Planner traverses the intersection without any collision.
As noted in the previous section, the no ego at-fault collisions metric does not penalize rear-end collisions. However, because EMoE-Planner’s interaction-oriented loss explicitly accounts for yielding and overtaking behaviors, rear-end interactions are also considered during optimization, enabling the model to effectively avoid such collisions.
Figure 8: Closed-loop simulation of left-turn scenario.
Figure 9: Closed-loop simulation of going straight scenario.
Figure 10: Closed-loop simulation of going straight within a junction.
Figure 11 presents the simulation results for a lane-splitting scenario, which falls under the ”others” category and involves challenges in lane choice and lane-change timing. DiffusionPlanner changes lanes too early, resulting in a collision at the 10-second mark. Pluto, on the other hand, drives too slowly and is rear-ended by a following vehicle at 12 seconds. While EMoE-Planner’s trajectory closely aligns with the ground truth, it comes dangerously close to a vehicle on the left at 14 seconds, indicating a potential collision risk. Nevertheless, EMoE-Planner achieves the best overall performance among the three planners in this scenario.
Figure 11: Closed-loop simulation of other scenarios.
The scenarios presented above are relatively common cases within each category. Figure 12 illustrates a more challenging right-turn scenario where the ego vehicle must navigate while multiple consecutive oncoming vehicles are executing left turns, creating a complex interaction zone with high collision risk. In this difficult setting, DiffusionPlanner again fails due to an excessively wide turning radius, resulting in a collision at 10 seconds. Conversely, PlutoPlanner adopts an overly conservative strategy, remaining almost entirely outside the interaction zone for the first 15 seconds and waiting for all oncoming traffic to clear before proceeding. EMoE-Planner successfully completes the right turn without collision. While its efficiency in this complex scenario is lower than that of the ground-truth trajectory, its overall behavior demonstrates a balanced approach: it remains cautious in the face of dense interactive traffic but continues to make progress without coming to a complete stop or moving excessively slowly, thereby maintaining a reasonable level of traffic throughput.
Figure 12: Closed-loop simulation of a difficult case.
Figure 13: Closed-loop simulation of a failure case.
Finally, Figure 13 presents a case that the current EMoE-Planner still struggles to handle: a head-on encounter on a bidirectional single-lane road where an oncoming vehicle partially intrudes into the ego’s lane. Insufficient proactive avoidance capability leads to a collision in this scenario. Although such interactions are theoretically covered by the interaction-oriented loss, they are extremely rare in the training data, so the model has not effectively learned a robust avoidance policy. Furthermore, in this particular setting, the available evasion space is very limited, and any evasive maneuver risks driving outside the drivable area, thus, the model refrains from taking action. The optimal strategy, as demonstrated by the ground-truth trajectory, would be to accelerate and pass through the conflict zone before the oncoming vehicle fully intrudes. This case highlights that the ability of EMoE-Planner to balance interactive safety and traffic efficiency in highly constrained, adversarial scenarios still requires further improvement, which will be a central focus of our future work.